# An Investigation of FP8 Across Accelerators for LLM Inference

### v2

@@
Abstract
The introduction of 8-bit floating-point (FP8) computation units in modern AI accelerators has generated significant interest in FP8-based large language model (LLM) inference. Unlike 16-bit floating-point formats, FP8 in deep learning requires a shared scaling factor. Additionally, while E4M3 and E5M2 are well-defined at the individual value level, their scaling and accumulation methods remain unspecified and vary across hardware and software implementations. As a result, FP8 behaves more like a quantization format than a standard numeric representation. In this work, we provide the first comprehensive analysis of FP8 computation and acceleration on two AI accelerators: the NVIDIA H100 and Intel Gaudi 2. Our findings highlight that the Gaudi 2, by leveraging FP8, achieves higher throughput-to-power efficiency during LLM inference, offering valuable insights into the practical implications of FP8 adoption for datacenter-scale LLM serving.
FP8, LLM, Quantization, GEMM, GEMV
1 Introduction
The FP8 format has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BF16 (Kalamkar et al., 2019), particularly at large batch sizes. With the introduction of the Intel Gaudi HPU and NVIDIA Hopper GPUs, FP8 is gaining traction as a key datatype for next-generation AI workloads (DeepSeek-AI, 2024).
While extreme low-bit quantization has been widely studied, its throughput remains inferior to conventional 16-bit floating-point formats on commercial hardware (Lin et al., 2024b). In contrast, FP8 has rapidly become the preferred format due to its compatibility with existing accelerators while preserving model accuracy. This makes it particularly well-suited for datacenter-scale LLM inference, where both high throughput and minimal accuracy degradation are critical. In some accelerators, FP8 can be twice as fast as BF16. Given the immense computational costs of training and deploying large language models (LLMs), FP8 has garnered significant attention for its potential efficiency gains.
Despite its advantages, FP8 utilization is constrained by the widening gap between computing speed and memory bandwidth. A key challenge is that, while LLM sizes continue to scale, their hidden dimensions grow at a much slower rate (e.g., Llama 8B has a hidden size of 4096, but Llama v3.3 70B only increases hidden size to 8192). Tensor parallelism (TP) further partitions matrices into even smaller fragments, making small matrix efficiency a critical performance factor.
Especially during the decode phase, Generalized Matrix-Vector (GEMV) and thin General Matrix-Matrix Multiply (GEMM) operations dominate, reducing Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021). With the growing popularity of “reasoning” models (DeepSeek-AI, 2025), the proportion of decoding in LLM inference workloads can be expected to continue growing.
To comprehensively address the challenges of LLM inference, we aim to quantify the various constraints that impact performance, including those introduced by FP8 quantization. To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
There is a notable lack of research providing comprehensive FP8 experimental results, particularly in studies analyzing hardware characteristics in LLM inference configurations. To address this gap, our work offers a detailed evaluation of FP8 performance across different hardware platforms, providing valuable insights for model developers on the trade-offs between model architecture and hardware design. To summarize our contributions:
Figure 1: TFLOPS (top row) and TFLOPS/Watt (bottom row) comparison between the Gaudi 2 and H100 using FP8 in the decode phase with batch size set to 64. We measure inference throughput using Optimum Habana v1.15.0 for the Gaudi 2 and TensorRT-LLM v0.16.0 for the H100. For short sequences, the Gaudi 2 has an advantage both in absolute throughput and in TFLOPS/Watt. At longer sequences, the H100 achieves higher TFLOPS due to its superior memory bandwidth, but is less energy efficient due to higher power consumption.
  1. 1.
    We analyze scaling factor configurations for FP8 formats and quantify key LLM inference constraints in datacenter environments, providing a unified framework to compare throughputs across different inference stages, model sizes, and sequence lengths.
  2. 2.
    We analyze the computational characteristics of LLM inference on the Gaudi 2 and H100 across both the prefill and decode phases. In the prefill phase, where compute performance is the key factor, H100’s throughput is 50% higher than Gaudi 2’s. However, since the decode phase dominates overall inference time, accelerator utilization becomes more critical, and Gaudi 2 achieves 50% higher throughput in this phase. This highlights the differing performance trade-offs between the two accelerators across LLM inference workloads.
2 Background
With the explosive growth of LLM sizes in recent years, numerous techniques specific to generative LLMs have emerged to reduce memory requirements and improve output speeds. In this work, we focus on the large-batch serving environment often found in data centers, where multiple requests for LLM inference run simultaneously under power constraints at both the rack and data center levels.
Various methods have been proposed to improve LLM serving efficiency, with quantization being one of the most widely studied. However, a key challenge in LLM quantization is the presence of activation outliers. Transformer models often contain large outlier values in their activations, which, if not properly addressed during quantization, can lead to substantial quality degradation (Sun et al., 2024a). To mitigate this issue, several quantization techniques have been developed, aiming to improve robustness and accuracy while maintaining computational efficiency.
2.1 Weight-only quantization
Due to the quantization difficulties posed by activation outliers, many studies have proposed several weight-only quantization (WoQ) methods(Lin et al., 2024a; Frantar et al., 2023; Lee et al., 2023), while retaining activations in BF16 or FP16 format. However, these approaches often suffer from low throughput at larger batch sizes. To enable computations, quantized weights must be converted into the same data type as activations (e.g. BF16), requiring online dequantization. This is handled by CUDA cores on NVIDIA GPUs and TPC cores on Gaudi HPUs, which operate an order of magnitude slower than their respective matrix multiplication units. Therefore, dequantization frequently becomes a bottleneck even at moderate batch sizes. Also, because GEMM computations are conducted at higher precisions, WoQ methods are fundamentally limited to having the same throughput as unquantized GEMMs at best.
Even when high throughput is achieved, kernel implementations often need to be re-implemented for different devices, including those from the same vendor. For instance, the Marlin kernel (Frantar et al., 2024), optimized for int4-weight and float16-activations on the A100, is suboptimal for the H100. As a result, the Machete kernel (Wilkinson, 2024) was developed specifically to leverage the new features of the Hopper architecture, employing different design trade-offs. These challenges arise due to variations in cache sizes, hardware features, and compute-to-memory bandwidth ratios, all of which necessitate different optimizations to achieve high performance. Consequently, adapting WoQ methods across different accelerators imposes significant implementation overhead and increases development costs.
2.2 INT8 quantization
The INT8 quantization format described in Xiao et al. (2024) is closely related to our work, requiring quantization of both weights and activations to leverage INT8 tensor cores to accelerate computations. While INT8 quantization has been widely adopted in fields such as computer vision, its limited dynamic range presents challenges in achieving full accuracy for LLM inference. FP8 thus has an advantage over INT8 as integer representations are inherently less robust to outliers compared to floating-point numbers (Shen et al., 2024; Kurtic et al., 2024).
2.3 FP8 quantization
The FP8 format can be categorized into two variants based on the allocation of bits to the exponent and mantissa: the E4M3 format and the E5M2 format. E4M3 offers higher precision compared to the E5M2 format, which instead has a higher dynamic range. To address the smaller dynamic range of FP8 compared to BF16, scaling factors are employed to adjust the inputs before casting to FP8. This ensures that the values fit within the FP8 range, maximizing the utilization of its limited dynamic range. Per-tensor scaling, which computes a single scaling factor for each matrix. For more granular control, per-token scaling can be applied to activations, while per-channel scaling is used for weights, enabling finer adjustments. Group-wise scaling factors (DeepSeek-AI, 2024) have also been proposed.
2.4 Scaling factor computation
A critical consideration for activation quantization is the overhead of converting activations to lower precision. While non-linear operations require 16-bit or even 32-bit precision, activations must be quantized online immediately before GEMM computations utilize low-precision compute units. This introduces non-trivial overhead, limiting the throughput gains from reduced precision. Activation quantization also requires estimating input statistics to determine appropriate scaling factors. A direct approach, measuring statistics for each activation before quantization, would necessitate two memory accesses: one for collecting statistics and another for quantization, incurring high latency. To mitigate this, many libraries use a calibration set to precompute scaling factors based on a representative sample of activations.
A key limitation of static scaling is that it typically enforces per-tensor scaling of activations, as token-wise variations are unknown in advance. For instance, altering the order of sequences during batched decoding would shift token positions. Moreover, using position-dependent scaling factors would require fixed input sizes. While per-channel scaling (Xiao et al., 2024) is possible, the same scaling factors are still applied to all tokens, potentially reducing accuracy.
In contrast, dynamic scaling is more efficient when applied separately to each token as measurement and quantization can be combined in a single memory access. It also enables group-wise quantization, as implemented in DeepSeek-AI (2024) with a group size of 128, improving output quality through finer-grained scaling factors.
A potential optimization for both scaling methods is to merge quantization with other memory-bound operations. In Pre-LN Transformers (Xiong et al., 2020), activation quantization can be fused with RMSNorm. Similarly, element-wise multiplication in SwiGLU (Shazeer, 2020) can also be fused. Frameworks such as TensorRT-LLM (NVIDIA, 2023) v0.16.0 have already implemented these optimizations for FP8 activation quantization. They have yet to be implemented for Gaudi HPUs as of the time of writing.
2.5 LLM evaluation
LLM evaluation has grown in importance as new LLMs compete for prominence. Traditionally, WikiText-2 perplexity (Merity et al., 2017) has been widely used due to its ease of measurement and relationship with the training loss. More recently, Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) has emerged as the de facto standard to assess LLM quality, evaluating knowledge in a wide range of domains. However, multiple-choice benchmarks may fail to capture the full impact of quantization. For instance, the Llama v3 405B model (Llama Team, 2024) applies max-absolute (maxabs) clipping when calculating scaling factors despite there being no effect in standard evaluation metrics because the model may generate corrupted outputs for high-perplexity tokens, such as dates, without maxabs clipping.
Moreover, evaluation should extend beyond pre-trained base models to instruction-tuned models that have undergone alignment. If quantization is applied after pre-training but before alignment, the alignment process must be quantization-aware, requiring each quantization algorithm to be integrated with each alignment method while maintaining alignment quality, significantly increasing complexity. For this reason, we conduct all evaluations on instruction-tuned models, where quantization effects are more representative of real-world deployment scenarios.
3 Matrix multiplication comparison
3.1 Differences in hardware capabilities
@@
(Accumulation precision) Hopper GPUs use a 14-bit accumulator for FP8 GEMMs (DeepSeek-AI, 2024), requiring casting to CUDA cores for higher precision. Software optimizations, such as applying FP32 accumulation to only one in four WGMMA instructions, reduce error but increase kernel complexity and remain less precise than full FP32 accumulation. In contrast, Gaudi HPUs always accumulate in FP32, ensuring higher numerical precision.
@@
(Power-of-2 scaling) Gaudi HPUs allow modification of floating-point exponent biases to accelerate scaling factor application. The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3 while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$. However, this feature is limited to GEMMs with per-tensor scaling factors.
@@
(Sparsity) NVIDIA GPUs support semi-structured sparsity acceleration, potentially doubling tensor core peak throughput. However, despite attempts to leverage sparsity in LLMs (Frantar & Alistarh, 2023; Sun et al., 2024b), dense GEMMs remain dominant due to accuracy loss and limited speedups. Gaudi HPUs do not support sparsity acceleration.
3.2 GEMM throughput measurements
As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance. While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices.
Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM. We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 24.12 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.19.0 image. Power consumption was recorded via the ”nvidia-smi” and ”hl-smi” utilities, respectively. Input casting overheads were excluded from the measurements. Additional results for GEMM throughputs under different conditions are provided in Tables 8 and 9. Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices. The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K. Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators. As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024). Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization.
This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes. Moreover, the lower power draw of the Gaudi 2 relative to its TDP suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
Table 1: Throughput and power measurements for square FP8 GEMMs with row-wise scaling. The ratio of measured TFLOPS and power draw relative to their peak values are included in parentheses. We include the TFLOPS/Watt ratio on the right. H100 has 1989.9 peak FP8 GEMM TFLOPS and 700W TDP. Gaudi 2 has 865 peak FP8 GEMM TFLOPS and 600W TDP.
Device (M,K,N) TFLOPS Power (W) TF/W
@@ 3.2 GEMM throughput measurements
4 Comparing end-to-end results
Table 2: The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration. The “Measured” column includes results for dynamic “FP8 (D)” scaling. Reference BF16 accuracies are included for both columns to control for evaluation condition differences.

### v3

# An Inquiry into Datacenter TCO for LLM Inference with FP8

@@
Abstract.
As large language models (LLMs) continue to scale, their inference demands present significant challenges, particularly due to the high power consumption of AI accelerators in datacenters. These facilities require specialized cooling and power management systems, substantially increasing the total cost of ownership (TCO) for cloud service providers (CSPs). In this work, we analyze the computational characteristics and constraints of LLM inference from a TCO perspective, focusing on two representative accelerators: the Gaudi 2 and NVIDIA H100. We present a generalizable framework that enables CSPs to compare and select AI accelerators according to diverse operational requirements. Using this model, we analyze the impact of FP8 precision and LLM inference workload characteristics as key factors influencing TCO. We investigate FP8 quantization, which is gaining adoption in LLM training, as a technique to improve inference throughput while maintaining cost efficiency. Furthermore, our analysis of LLM inference workloads reveals that performance on thin GEMMs, which dominate the decode phase, can have a greater impact than theoretical hardware peak performance. By studying the interaction between power consumption, quantization strategies, and hardware architecture, we offer insights that support informed deployment decisions and guide future accelerator designs to improve the TCO of LLM inference.
LLM inference, TCO analysis, FP8 quantization
conference: Arxiv; 2025; Arxiv
1. Introduction
With the rapid advance of large language models (LLMs), AI datacenters have emerged as a critical infrastructure for supporting scalable and efficient inference workloads. The recent emergence of ”reasoning” AI models for both proprietary models such as OpenAI’s o1 (OpenAI, 2024) and Claude 3.7 (Anthropic, 2025), as well as open-source models such as DeepSeek R1 (DeepSeek-AI, 2025), has solidified the expectation that test-time scaling of LLMs will become a central workload for cloud service providers (CSPs). As LLM inference requests surge in both quantity and complexity, the volume of output tokens per request is also expected to grow substantially. This trend places significant pressure on datacenter infrastructure, making the selection of appropriate hardware a critical factor for sustaining performance and service-level efficiency. In this context, competition among AI accelerator manufacturers has intensified, with each aiming to position themselves to take advantage of the market.
Hardware selection has a direct impact on the total cost of ownership (TCO), which in turn influences the overall profitability and operational viability of AI datacenters. This necessitates a detailed examination of the various components that impact TCO, ultimately aiming to identify effective strategies for minimizing it. To provide practical guidance for CSPs, we construct an analytical model to compare hardware options, taking into account both facility costs and target performance levels. Although extensive research has explored the design of advanced AI accelerators (Micikevicius et al., 2022; Lee et al., 2025) and LLM quantization techniques (Lin et al., 2024; Ashkboos et al., 2024; Xiao et al., 2024), there remains a lack of rigorous investigation into how these developments impact the TCO for datacenter-scale LLM inference. Building upon our model, we analyze several key factors that influence TCO and discuss potential approaches for its optimization.
First, we analyze the impact of data precision in LLMs, with a particular focus on FP8 quantization, which has seen increasing adoption in recent state-of-the-art models such as DeepSeek v3 (DeepSeek-AI, 2024) and Llama 4 (Meta, 2025). Quantization offers the additional benefit that modern AI accelerators often include dedicated hardware units designed to accelerate low-precision computation, thereby improving throughput. In particular, FP8 has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BrainFloat16 (BF16) (Kalamkar et al., 2019), especially for large batch sizes. With the adoption of FP8 by next-generation accelerators such as Intel’s Gaudi HPU and NVIDIA’s Hopper GPU, this format is gaining traction in cutting-edge LLMs. In this context, we explore how different FP8 quantization strategies impact both accuracy and throughput, examining their effects at two levels: matrix multiplication operations and end-to-end LLM inference.
Second, we investigate the behavior of LLM inference workloads when executed with FP8 quantization, aiming to understand its impact on throughput and resource utilization. To leverage quantization in a cost-effective manner, it is essential to understand the computational characteristics of LLM inference, which differ significantly from those of training. Even within the inference process, throughput varies across two distinct phases: the prefill phase and the decode phase. In particular, the decoding phase is dominated by Generalized Matrix-Vector (GEMV) operations and thin General Matrix-Matrix Multiplication (GEMM), both of which exhibit low computational intensity (CI) and lead to reduced Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021). Furthermore, performance varies significantly across hardware platforms, depending on factors such as memory interface bandwidth and matrix computation unit throughput. Therefore, we conduct an in-depth analysis of the characteristics of the H100 and Gaudi2 in the context of LLM inference workloads. We further identify the performance limitations of each accelerator by examining their inherent architectural bottlenecks. To summarize our contributions:
  1. (1)
    We present a scalable TCO estimation model that serves as a generalizable framework for selecting optimal hardware configurations tailored to the needs of CSPs. While previous studies (Patel et al., 2024; Griggs et al., 2024) have analyzed throughput-cost trade-offs, they are often limited to specific deployment scenarios. In contrast, our model accommodates a wide range of cost structures and performance requirements, enabling flexible evaluation across different CSP contexts. Moreover, the model offers intuitive guidance for TCO optimization from a throughput-centric perspective.
  2. (2)
    Using this model, we investigate two key factors influencing TCO: FP8 numerical precision and LLM inference workload characteristics. Despite FP8’s growing popularity as a low-precision format that offers favorable trade-offs between throughput and memory efficiency, prior work has largely overlooked its empirical evaluation, particularly in the context of hardware-specific behavior during LLM inference. To address this gap, we conduct a comprehensive analysis of FP8 performance across different accelerator platforms. In addition, we provide a focused analysis of LLM decoding throughput that goes beyond a simple inspection of hardware specifications. Under FP8-quantized conditions, we further explore the performance characteristics and bottlenecks of each device within LLM workloads, offering insights that inform both hardware design and deployment strategies.
2. Exploring the TCO-throughput trade-off
People who buy and sell chips think about the cost of chips. People who operate data centers think about the cost of operations. … Our TCO is so good that even when the competitor’s chips are free, it’s not cheap enough.
Jensen Huang, Founder & CEO of NVIDIA
2.1. Understanding the TCO model
Figure 1. TCO comparison table for fixed traffic. For any desired task, given the throughput ratio between accelerators A and B $R_{Th}$ and server price ratio $R_{SC}$, the table shows the TCO of accelerator A as a proportion of the TCO of accelerator B, with $C_{S}=C_{I}$ assumed for clarity. A is therefore more cost-efficient than B when the obtained value is less than 1. The lower the value, the greater the cost-effectiveness of A.
When calculating the costs involved in running AI accelerators for datacenter operators, one must note that the purchase price of the chips themselves, while significant, constitute only part of the Total Cost of Ownership (TCO). For example, the purchase cost of the networking, cooling, and power infrastructure necessary to maintain the processors is frequently as large as the purchase cost of the chips and may even exceed it (Cui et al., 2017). As AI inference workloads constitute an ever greater portion of datacenter operations, achieving low cost and high efficiency requires careful planning. The choice of computing systems impacts both the throughput and the power consumption, both of which have a substantial influence on TCO.
Consider a scenario in which a CSP must choose between accelerators for a datacenter. In this scenario, two alternatives exist: system A, which is lighter-weight, delivers lower peak specifications, and has a lower purchase price, and system B, which provides substantially higher peak specifications but comes with a higher per accelerator purchase price. To identify which hardware option yields higher profitability, we must construct a model to compare TCO between the two hardware options.
When purchasing a server, it contains not only the chip but also networking, storage, and other associated components required for the server to function. Also, mulltiple chips may be located in a single server, whose costs must also be accounted for. In the following discussion, we denote $ServerCost$ as the term that combines the purchase costs of the multiple accelerator chips, CPUs, storage, and networking devices required to constitute a operational server.
$InfraCost$ indicates the infrastructure costs associated with operating the datacenter, including cooling and power provision, divided by the number of accelerators. We simplify the analysis by including operating expenses (OpEx), such as electricity consumption, in $InfraCost$. To understand the infrastructure cost, note that while the cost of electricity increases almost linearly with regard to the number of processors, the cost of the associated infrastructure, such as power racks and cooling equipment, is nearly constant. Because datacenter racks can only provide a limited amount of power, the per-chip cost of infrastructure is inversely proportional to the number of servers that can fit in a single rack. Therefore, the benefits of lower power consumption are twofold: the electricity consumption goes down and the number of processors per rack increases, spreading fixed infrastructure costs among a greater number of servers. For most datacenters, the cost of electricity per se is outweighed the cost of the rack and other equipment, even when amortized over several years, leading $InfraCost$ to be most heavily affected by the number of chips per rack.
We denote $R_{SC}$, $R_{IC}$, and $R_{Th}$ as the relative server costs, per server infrastructure costs, and throughputs between a single server of computing systems A and B, respectively, and define them as follows:
$R_{SC}=\frac{ServerCost_{A}}{ServerCost_{B}},\quad R_{IC}=\frac{InfraCost_{A}}{InfraCost_{B}},\quad R_{Th}=\frac{Throughput_{A}}{Throughput_{B}}$
Regardless of the devices in use, the amount of traffic remains constant. We therefore make an iso-traffic assumption and use $N$ to represent the number of B servers required to process the expected traffic. We use $C_{S}=ServerCost_{B}$ and $C_{I}=InfraCost_{B}$ to represent the cost components attributed to the server and the infrastructure cost per server when using system B. Making the assumption that the cost of infrastructure is the same for systems A and B, i.e., $R_{IC}=1$, derive the following equation:
(1) $\frac{TCO_{A}}{TCO_{B}}=\frac{C_{S}\times R_{SC}\times\frac{N}{R_{Th}}+C_{I}\times\frac{N}{R_{Th}}}{C_{S}\times N+C_{I}\times N}$
The numerator captures the TCO incurred when using the lower-performance system A to meet the same throughput target as system B, while the denominator reflects the baseline TCO of using system B for a given amount of traffic. As total traffic remains constant regardless of the underlying device, purchasing system A will require $R_{Th}/N$ nodes, leading to possibly increased infrastructure costs, depending on the throughput ratio $R_{Th}$. At the same time, its lower per-chip cost, reflected in the server price ratio $R_{SC}$, could result in a lower TCO.
Based on this model, Figure 1 illustrates the TCO comparison between two computing systems, expressed as an $A/B$ ratio. The x-axis represents the server price ratio, and the y-axis represents the per server throughput ratio of the two systems for the desired task. Since exact costs are often confidential in real-world data centers, expressing TCO in relative ratios allows for flexible adaptation to various operational contexts. In deriving the values in Figure 1, we assumed $C_{S}=C_{I}$, assuming that server and infrastructure costs contribute equally (i.e., 50% each) to the overall TCO, thereby allowing these terms to cancel out across the numerator and denominator. We also assume that both systems consume the same power per server and that the same number of servers fit in a single rack, i.e., $R_{IC}=1$. In practice, however, it may be appropriate to substitute variables to reflect each specific deployment scenario. The values in Figure 1 are an example of the TCO ratio between two hardware options. The lower the value, the less expensive, and hence more cost-effective, system A becomes and the more appealing it is to adopt. Conversely, a ratio above 1.0 suggests that system B is preferable from a TCO perspective.
2.2. TCO factor analysis
The TCO analysis presented so far offers a flexible and adaptable methodology for evaluating system options across diverse data center environments. In the implemented TCO model, both the server price ratio and throughput ratio can vary significantly depending on factors such as server acquisition costs, accuracy constraints, and the specific tasks being executed. The cost of processors may change depending on negotiations with manufacturers and are generally confidential.
A key contribution of the analysis is that it can incorporate throughput ratios of any desired task, e.g., LLM decoding as opposed to LLM training or prefill, allowing users to take the intricacies of hardware-software co-optimization into account. For instance, the service may require that only models that pass a quality assurance (QA) review are allowed, and that the model meets specific latency requirements when deployed. All of these factors can be incorporated into the TCO model to provide insights. For example, the prefill and decode phases of LLM inference exhibit markedly different throughput profiles. As shown in Patel et al. (2024), it is possible to use different chips and configurations for the prefill and decode to optimize TCO as their relative throughputs for the prefill and decode tasks are different. When these different throughputs are compared using our TCO model, it becomes easier to grasp why it is profitable to utilize different hardware and configurations for different inference phases.
To build an efficient datacenter environment, CSPs must consider complex trade-offs between functionality, performance, and cost. Quantization has emerged as a viable solution to overcome some of these limitations by reducing memory usage. Leveraging the inherent error resilience of LLMs enables reduced bit-width representations of model parameters while maintaining comparable output quality. Consequently, incorporating these lower-precision modes into TCO analysis becomes essential. In the following section, we explore FP8 quantization as a particularly important technique.
3. FP8 for high-throughput LLM inference
In this section, we provide an overview of FP8 as a numerical format, analyze how it is implemented across hardware backends, and evaluate its performance in GEMM operations, one of the core building blocks of LLM inference.
3.1. Motivation for FP8 precision
While BF16 precision is commonly used in recent LLMs, there is substantial opportunity to boost throughput and to minimize TCO by operating at a lower precision level, FP8, within acceptable bounds of accuracy degradation. Especially with the emergence of next-generation foundation models such as DeepSeek v3 and LLaMA 4 whose parameters were trained in FP8 precision, the FP8 GEMM has become increasingly central to achieving high throughput at scale. Moreover, even for models originally trained using BF16, post-training quantization (PTQ) techniques now enable effective conversion to FP8 without significant accuracy degradation. However, the effectiveness of low-precision computation is highly dependent on the underlying hardware architecture. As precision decreases, achieving full hardware utilization becomes more challenging, directly affecting throughput. Therefore, hardware-specific characteristics are critical in determining actual performance at lower precision levels. In this section, we explore the characteristics and implications of Llama 4 style FP8 precision, analyzing its computational behavior and potential for performance optimization.
3.2. Differences in hardware capabilities
@@
(Accumulation precision) Hopper GPUs use a 14-bit accumulator for FP8 GEMMs (DeepSeek-AI, 2024), requiring casting to CUDA cores for higher precision. Software optimizations, such as applying FP32 accumulation to only one in four WGMMA instructions, reduce error but increase kernel complexity and remain less precise than full FP32 accumulation. In contrast, Gaudi HPUs always accumulate in FP32 (Lee et al., 2025), ensuring higher numerical precision.
@@
(Power-of-2 scaling) Gaudi HPUs allow the modification of floating-point exponent biases to accelerate scaling factor application. The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3, while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$. However, this feature is limited to GEMMs with per-tensor scaling factors.
@@
(Sparsity) NVIDIA GPUs support semi-structured sparsity acceleration, potentially doubling tensor core peak throughput. However, despite attempts to leverage sparsity in LLMs (Frantar and Alistarh, 2023; Sun et al., 2024), dense GEMMs remain dominant due to accuracy loss and limited speedups. Gaudi HPUs do not support sparsity acceleration.
Table 1. Throughput and power measurements for square FP8 GEMMs with row-wise scaling. Ratios of measured TFLOPS and power to their peak values are shown in parentheses. H100 has peak FP8 throughput of 1989.9 TFLOPS with a TDP of 700W, and Gaudi 2 has 865 TFLOPS with a TDP of 600W.
Device (M,K,N) TFLOPS Power (W) TFLOPS/W
@@
3.3. GEMM throughput measurements
As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance. While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices. Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM. We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 25.03 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.12.0 image. Power consumption was recorded via the “nvidia-smi” and “hl-smi” utilities, respectively. Input casting overheads were excluded from the measurements. Additional results for GEMM throughputs under different conditions are provided in Tables 2 and 3. Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices. The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K. Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators. As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024). Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization. This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes. Moreover, the lower power draw of the Gaudi 2 relative to its TDP, as shown in Table 1, suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
Table 2. Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses. Hardware acceleration is only available for per-tensor scaling.
Type Size Per-Row Per-Tensor HW Accel.
E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
Table 3. H100 throughput (TFLOPS) for scaled FP8 GEMMs on square matrices of the given size, excluding quantization overhead. Ratios relative to peak FP8 throughput are shown in parentheses. Only E4M3 results are reported, as Python functions for E5M2 GEMM are unavailable for GPUs.
Accum. Size Per-Row Per-Tensor
FP32 1K 217 (11.0%) 186   (9.4%)
2K 299 (15.1%) 840 (42.4%)
4K 362 (18.3%) 1099 (55.5%)
8K 396 (20.0%) 1300 (65.7%)
Fast 1K 237 (12.0%) 147   (7.4%)
2K 810 (40.9%) 896 (45.3%)
4K 1136 (57.4%) 1205 (60.9%)
8K 1123 (56.8%) 1388 (70.1%)
4. FP8 configuration impact on accuracy
Table 4. The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration. The “Measured” column includes results for dynamic “FP8 (D)” scaling. BF16 accuracies are included for both columns to control for evaluation condition differences.